OpenStudy (anonymous):

Find the area of a circle that has a diameter of 11 inches. Round your answer to the nearest hundredth.

OpenStudy (pinksapphire):

A=3.14*5.5^2

OpenStudy (anonymous):

π D = 11 D=11/π D = 3.501408 D/2 = r r = 1.7507 π r^2 = 9.63

OpenStudy (pinksapphire):

95.033

OpenStudy (anonymous):

But why would it not be 95.03

OpenStudy (pinksapphire):

Yes it would be 95.03...Sorry about that :(

OpenStudy (anonymous):

So it is 95.03 not 9.63

OpenStudy (pinksapphire):

Yes.

OpenStudy (anonymous):

no she made a mistake, to find the diameter you need to multiply π by D to get diameter instead she multiplied D * 2 to find diameter which is whrong\

OpenStudy (pinksapphire):

This is the formula to find the area of a circle: \[A =πr^2\]

OpenStudy (pinksapphire):

5.5 is the radius.

OpenStudy (anonymous):

nope the radius is half the diameter And the diameter is Circumference divided by π

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

oh wait sorry I was reading two problems at the same time, PinkSohire is actually right

OpenStudy (pinksapphire):

I know that...the diameter is 11 so to get the radius, you have to divide by 2.

OpenStudy (anonymous):

yeah sorry I thought it said circumfernce

OpenStudy (pinksapphire):

Oh okay :)

OpenStudy (anonymous):

So 95.03 is correct

OpenStudy (pinksapphire):

Yes.